OpenStudy (anonymous):

If f(x) = x - 2, for what value of x does f (3x) = f(x+1)?

OpenStudy (anonymous):

\[f(x)=x-2\] so \[f(3x)=3x-2\] similarly \(f(x+1)=x+1-2=x-1\) set \[3x-2=x-1\] and solve for \(x\)

OpenStudy (shamim):

given that \[f(x)=x-2\]

OpenStudy (shamim):

so\[f(3x)=3x-2\]

OpenStudy (shamim):

and\[f(x+1)=x+1-2\]

OpenStudy (shamim):

try to complete the rest

OpenStudy (anonymous):

Hint: f(3x) = 3x - 2 and f(x + 1) = x + 1 - 2

OpenStudy (anteater):

f(x) = x - 2, so f(3x) = 3x - 2 and f(x+ 1) = x + 1 - 2 or x -1 f(3x) = f(x + 1) --> 3x - 2 = x - 1 --> 2x - 2 = -1 --> 2x = 1 --> x = 1/2 Verify: f(3X1/2) = f(3/2) = 3/2 - 2 = 3/2 - 4/2 = -1/2 f(1/2 + 1) = f(3/2) = -1/2